### v1

@@ 6.5 Variable Selection Performance under Highly Correlated Predictors
Table 3: Variable selection performance under highly correlated predictors. Avg. total: average number of selected variables; Avg. false: average number of falsely selected noise variables; Noise sel. rate: average per-noise-variable selection probability.
@@
7 Conclusion
This work proposes a unified DC modeling framework, termed DCEN, by integrating the nonconvex $\ell_{1}-\alpha\ell_{2}$ difference term with the convex $\ell_{2}^{2}$ regularization. This formulation achieves an effective balance between strong sparsity promotion and solution stability, alleviating the bias of $\ell_{1}$ regularization while preserving robustness under highly correlated variables. With continuously tunable parameters, DCEN provides a unified modeling paradigm that encompasses and generalizes several representative models, including Lasso, Elastic Net, and the nonconvex $\ell_{1}-\alpha\ell_{2}$ model.
From a theoretical and algorithmic perspective, sufficient conditions for exact and stable recovery of CDCEN are established under the RIP framework. Closed-form expressions of the proximal operator associated with the DCEN regularizer are derived, enabling the development of efficient optimization algorithms based on DCA and ADMM through DC decomposition and variable splitting. Moreover, within the KŁ analytical framework, the global convergence of DCA is rigorously proved, and its linear convergence rate is further characterized.
Furthermore, DCEN is extended to image processing applications by incorporating total variation regularization, leading to DCEN-TV, which is efficiently solved via a split Bregman method. Extensive numerical experiments on high-dimensional variable selection with strong correlations and MRI image reconstruction demonstrate that the proposed approach consistently outperforms state-of-the-art methods in terms of recovery accuracy, robustness, and noise resilience. Future work will focus on strengthening theoretical recovery guarantees, and extending DCEN to broader applications, including low-rank matrix recovery and multitask learning.

### v2

@@ VII-E Variable Selection under Highly Correlated Predictors
TABLE III: Variable selection performance under highly correlated predictors. Avg. total: average number of selected variables; Avg. false: average number of falsely selected noise variables; Noise sel. rate: average per-noise-variable selection probability.
@@
VIII Conclusion
This work proposes DCEN, a unified Difference-of-Convex modeling framework that integrates the nonconvex $\ell_{1}-\alpha\ell_{2}$ penalty with $\ell_{2}^{2}$ regularization. By balancing sparsity promotion with solution stability, DCEN generalizes several representative models—including LASSO, Elastic Net, and the nonconvex $\ell_{1}-\alpha\ell_{2}$ model—effectively alleviating $\ell_{1}$-induced bias while maintaining robustness under high variable correlation.
Theoretically, we establish a comprehensive statistical consistency framework for DCEN. We derive oracle inequalities and provide rigorous recovery bounds that characterize the estimator’s behavior in high-dimensional regimes. Furthermore, we establish sufficient conditions for exact and stable recovery under RIP. Algorithmically, we derive closed-form expressions for the proximal operator of the DCEN regularizer, enabling efficient optimization via DCA and ADMM. Within the Kurdyka-Łojasiewicz (KŁ) framework, we prove the global convergence of DCA and characterize its linear convergence rate.
Extending the framework to image processing, we develop DCEN-TV using total variation regularization, solved via the split Bregman method. Extensive numerical experiments in high-dimensional variable selection and MRI reconstruction demonstrate that DCEN consistently outperforms state-of-the-art methods in accuracy, robustness, and noise resilience. Future research will explore extensions to low-rank matrix recovery and multitask learning.